OpenStudy (anonymous):

Solve: 5 over quantity of 4 plus square root of 5

OpenStudy (mertsj):

Multiply numerator and denominator by the conjugate of 4+sqrt5

OpenStudy (mertsj):

\[\frac{5}{(4+\sqrt{5})}\times\frac{(4-\sqrt{5})}{(4-\sqrt{5})}\]

OpenStudy (mertsj):

Can you finish that?

OpenStudy (amistre64):

thats a viable option, but how is that a solution to an expression?

OpenStudy (anonymous):

so its 20 - sqrt of 5 over 11?

OpenStudy (mertsj):

You getting warm. Multiply both terms of the binomial by 5 in the numerator

OpenStudy (mertsj):

\[5(4-\sqrt{5})=20-5\sqrt{5}\]

OpenStudy (anonymous):

and then that over 11?

OpenStudy (mertsj):

Yes.